# Reconstruction-based Multi-Normal Prototypes Learning for Weakly Supervised Anomaly Detection

### v1

@@ 3.4 Multi-Normal Prototypes Learning Module
Table 2: Performance comparison of different methods. The upper section presents AUC-PR (primary metric) and the lower section shows AUC-ROC. The best results are highlighted in bold and the second-best results are underlined.
3.5 Unified Anomaly Scoring Module
To detect anomalies with consideration of both the reconstruction error $e_{i}=\mathit{l}(x_{i},\hat{x}_{i})$ and the multi-normal prototype information, we design a unified anomaly scoring module. Inspired by the residual network (He et al. 2016), we also take the latent representation $z_{i}$ of the given sample $x_{i}$ as the input. In summary, the concatenated vector forms the input to the unified anomaly score evaluator $\varphi$ to get an anomaly score:
$score_{i}=\varphi(\text{cat}(e_{i},z_{i},s_{i});\Theta_{\varphi})$ (11)
3.6 Training and Inference
Training.
Our model is trained in an end-to-end manner using mini-batches consisting of $b_{u}$ unlabeled samples and $b_{a}$ labeled anomalies. For each mini-batch, we compute three losses: $L_{recon}$, $L_{np}$, and $L_{score}$, as defined by equation (3), (10), and (12), respectively. To effectively balance these three losses dynamically, we utilize a technique known as Dynamic Averaging, inspired by (Liu, Johns, and Davison 2018). The total loss is calculated as follows:
$\mathcal{L_{\text{total}}}=w_{1}\mathcal{L_{\text{{recon}}}}+w_{2}\mathcal{L_{\text{np}}}+w_{3}\mathcal{L_{\text{score}}}$ (13)
where $w_{1}$, $w_{2}$, and $w_{3}$ are updated per epoch based on the rate of descent of the corresponding losses.
Inference.
4 Experiments
4.1 Experimental Setup
Datasets
This section outlines the datasets used in our experiments, covering a broad range of scenarios across domains like network security, Internet of Thing (IoT) security, financial fraud detection, and others. All 15 datasets are summarized in Table 1, providing robust benchmarks for evaluating anomaly detection methods with their diverse scenarios and varying anomaly ratios.
Baseline Methods
This section outlines the algorithms that are most comparable to our proposed method. These algorithms are grouped into two categories based on their supervision level: 1) Weakly supervised algorithms, such as RoSAS (Xu et al. 2023b), PReNet (Pang et al. 2023), DevNet (Pang, Shen, and van den Hengel 2019), DeepSAD (Ruff et al. 2020), and FeaWAD (Zhou et al. 2021), leverage limited labeled data alongside a larger set of unlabeled data, offering a direct comparison to our approach; 2) Unsupervised algorithms like DeepIForest (Xu et al. 2023a), DeepSVDD (Ruff et al. 2018a), and iForest (Liu, Ting, and Zhou 2012), operate without any labeled data, offering a baseline to assess the benefits of incorporating even minimal labeled data in anomaly detection.
Evaluation Metrics
@@ 4.1 Experimental Setup
We use the Area Under the Precision-Recall Curve (AUC-PR) as the primary evaluation metric, complemented by the Area Under the Receiver Operating Characteristic Curve (AUC-ROC). AUC-PR is ideal for imbalanced datasets, emphasizing correct anomaly detection while minimizing false positives and negatives. Both metrics range from 0 to 1, with higher values indicating better performance. Results are averaged over 5 runs, and the paired Wilcoxon signed-rank test (Woolson 2008) is used to validate the significance of our findings.
Implementation Details
@@
We scale each feature to [0, 1] via min-max normalization. We use the Adam optimizer with a learning rate of $5\times 10^{-3}$ and weight decay of $5\times 10^{-4}$. The hidden representation dimension $H$ is set to 8, and the hyperparameters are configured as follows: batch sizes $b_{u}=128$, $b_{a}=32$; margin $m_{1}=0.02$; Student’s t-distribution degree of freedom $\alpha=1$; and similarity scaling parameter $\beta=1$. The number of normal prototypes is determined via clustering analysis on the unlabeled samples. Implementation of the compared methods is based on the DeepOD framework (Xu et al. 2023a).
4.2 Performance Comparison
@@
In our experiments, we evaluate the performance of our proposed anomaly detection method across multiple datasets with a fixed labeled anomaly ratio of 1.0% for each dataset. The labeled anomaly ratio is defined as the proportion of anomalies in the training dataset that are labeled. Table 2 shows the performance of different methods on the fifteen benchmark datasets. Our method consistently outperformed other approaches, achieving the highest average scores in both AUC-PR and AUC-ROC metrics, with statistically significant improvements with over 99.9% confidence. Specifically, in terms of AUC-PR, our method achieved the best results on 13 of the 15 datasets, and was the second-best on the remaining two, closely approaching the top results. The average AUC-PR of our method was 0.642, which is significantly higher than other methods. Similarly, for the AUC-ROC metric, our method achieved an average of 0.886, surpassing all other methods. The statistical validation further confirmed that the improvements brought by our method are significant, providing robust and reliable anomaly detection performance in weakly supervised settings. Besides, the weakly supervised algorithms, such as our method, RoSAS, and DevNet, perform significantly better than all unsupervised algorithms on almost all datasets. This confirms that with the appropriate use of a few labeled anomaly samples, we can get better anomaly detection performance.
@@
Table 3: Detection of unseen anomalies, with performance measured by AUC-PR. The ”Seen” indicates the anomaly type included in the training and validation sets, while the ”Unseen” indicates the anomaly type only included in the test set.
4.3 Effects of More or Less Labeled Anomalies
@@
This experiment aims to examine the effects of Labeled Anomaly Ratios on the model’s effectiveness in detecting anomalies, helping us understand the trade-off between supervision level and detection performance. From Figure 3, it is evident that our method consistently achieves the best performance across nearly all Labeled Anomaly Ratios. Specifically, as the Labeled Anomaly Ratios decrease, our method’s performance degrades more slowly and maintains relatively high detection performance, while other methods exhibit a significant decline in most datasets. On the other hand, from the perspective of anomaly contamination, as the proportion of labeled anomalies decreases, the proportion of unlabeled anomalies contaminating the unlabeled data increases. Figure 3 shows that, unlike other methods, our method’s detection performance declines more slowly as the anomaly contamination ratio increases, maintaining a relatively good detection performance. This demonstrates the robustness of our model against anomaly contamination.
Method Average AUC-PR P-value
w/o $\mathcal{L_{\text{recon}}}$ 0.622 ± 0.038 0.0008
w/o $\mathcal{L_{\text{np}}}$ 0.618 ± 0.022 0.0038
w/o $w_{i}$ 0.602 ± 0.042 $<$0.0001
w/o Decoder 0.611 ± 0.052 $<$0.0001
w/o MNP 0.566 ± 0.021 $<$0.0001
Full Model 0.642 ± 0.025 -
Table 4: Results of ablation study.
4.4 Detection Performance on Unseen Anomalies
@@
A crucial aspect of an effective anomaly detection system is its ability to generalize from known anomalies to detect previously unseen ones. This capability is essential for maintaining robustness in dynamic environments where new types of anomalies may emerge. To evaluate our method’s performance on unseen anomalies, we constructed several experimental datasets using the UNSW-NB15 dataset, which contains four types of anomalies: DoS, Rec, Ana, and Bac. In each experiment, two types of anomalies were selected: one used exclusively in the test set (unseen) and the other included in the training and validation sets (seen). This setup resulted in 12 different experimental datasets.
@@
The results are presented in Table 3, showing the AUC-PR scores for different methods across various combinations of seen and unseen anomaly types. Our method achieved the highest average AUC-PR of 0.765, significantly outperforming other methods by more than 20%. The statistical confidence level for this improvement exceeds 99.0%. This experiment demonstrates that our method can effectively generalize from known anomalies to detect previously unseen ones, maintaining strong detection performance and robustness in dynamic environments.
4.5 Ablation Study
We conducted an ablation study to assess the impact of each component in our model, as shown in Table 4. The results demonstrate that the full model achieves the highest AUC-PR of 0.642. Specifically, excluding the reconstruction loss or the normal prototype loss, deleting weight $w_{i}$, and removing the Decoder included reconstruction error all resulted in decreased AUC-PR values. Notably, the most significant performance degradation occurred when the multi-normal prototypes (MNP) were excluded, with the AUC-PR dropping to 0.566, highlighting the essential role of multi-prototype learning in our model.
@@ 4.5 Ablation Study
4.6 Sensitivity Analysis
@@
In this section, we investigate the effects of varying the number of normal prototypes. As shown in Figure 4 (illustrating four representative datasets), most datasets achieve better results with multiple prototypes, confirming our hypothesis that normal data is multi-categorical. In the Rec dataset, prototype count has little effect, indicating lower sensitivity. However, in the C&C and Census datasets, the choice of prototype number is crucial—too many or too few prototypes can significantly degrade performance. In the Bank dataset, although performance fluctuates with the number of prototypes, it consistently outperforms using a single prototype.
5 Conclusions

### v2

@@ IV-A2 Baseline Methods
IV-A3 Evaluation Metrics
@@
The same as previous studies [17, 15, 16], we primarily use the Area Under the Precision-Recall Curve (AUC-PR) as our main evaluation metric, supplemented by the Area Under the Receiver Operating Characteristic Curve (AUC-ROC). The AUC-PR is particularly effective for imbalanced datasets, emphasizing the model’s ability to correctly identify anomalies (true positives) while minimizing false positives and false negatives, making it ideal for anomaly detection where anomalies are rare. The AUC-ROC provides an aggregate measure of the model’s performance across all classification thresholds, illustrating its ability to distinguish between positive and negative classes. These two metrics both range from 0 to 1 and higher values indicate better performance. To ensure robust and reliable results, we report the average performance over 5 independent runs and employ the paired Wilcoxon signed-rank test[48] to statistically validate the significance of our findings.
IV-A4 Implementation Details
@@
We use the Adam optimizer with a learning rate of $5\times 10^{-3}$ and weight decay of $5\times 10^{-4}$. The latent representation dimension $H$ is set to 8, and the hyperparameters are configured as follows: batch sizes $b_{u}=128$, $b_{a}=32$; margin $m_{1}=0.02$; Student’s t-distribution degree of freedom $\alpha=1$; and similarity scaling parameter $\beta=1$. The number of normal prototypes k is determined via clustering analysis on the unlabeled samples. The implementation of the compared methods is primarily based on the DeepOD framework[46] and scikit-learn (sklearn).
IV-B Performance Comparison
The same as in [1], we evaluate the performance of our proposed anomaly detection method across multiple datasets with a fixed labeled anomaly ratio of 1.0% for each dataset. The labeled anomaly ratio is defined as the proportion of anomalies in the training dataset that are labeled. This means that in practical anomaly detection scenarios, we only need to label anomalies included in 1.0% of the training data, ensuring that the weakly supervised anomaly detection setup does not affect the proportion of anomalies in the training set regardless of the inherent anomaly ratio of the dataset.
@@ IV-B Performance Comparison
Table II shows the performance of different methods on the fifteen benchmark datasets. Our method consistently outperformed other approaches, achieving the highest average scores in both AUC-PR and AUC-ROC metrics, with statistically significant improvements with over 99.9% confidence. Specifically, in terms of AUC-PR, our method achieved the best results on 13 of the 15 datasets, and was the second-best on the remaining two, closely approaching the top results. The average AUC-PR of our method was 0.642, which is significantly higher than other methods. Similarly, for the AUC-ROC metric, our method achieved an average of 0.886, surpassing all other methods. The statistical validation further confirmed that the improvements brought by our method are significant, providing robust and reliable anomaly detection performance in weakly supervised settings.
Besides, the weakly supervised algorithms, such as our method, RoSAS, and DevNet, perform significantly better than all unsupervised algorithms on almost all datasets. This confirms that with the appropriate use of a few labeled anomaly samples, we can get better anomaly detection performance.
@@
TABLE III: Detection of unseen anomalies, with performance measured by AUC-PR. The ”Seen” indicates the anomaly type included in the training and validation sets, while the ”Unseen” indicates the anomaly type only included in the test set.
@@
IV-C Effects of More or Less Labeled Anomalies
This experiment aims to examine the effects of Labeled Anomaly Ratios on the model’s effectiveness in detecting anomalies, helping us understand the trade-off between supervision level and detection performance. The Labeled Anomaly Ratios refer to the proportion of labeled anomaly samples within the dataset. A higher ratio implies more known information about anomalies, bringing the learning process closer to a supervised setting and potentially improving the model’s anomaly detection capabilities, albeit at a higher labeling cost. Conversely, a lower ratio indicates less labeled information and lower labeling costs but challenges the model’s detection abilities due to reduced guidance. Because AUC-ROC and AUC-PR show almost consistent results, we only present AUC-PR results here. Additionally, due to space limitations, we only showcase a subset of datasets. In Fig. 3, the $x$ axis represents Labeled Anomaly Ratios ranging from 0.5% to 20%, and the $y$ axis represents the AUC-PR values for each dataset across all methods (except for unsupervised methods, as they are not affected by the labeled anomaly ratios).
@@ IV-C Effects of More or Less Labeled Anomalies
From Fig. 3, it is evident that our method consistently achieves the best performance across nearly all Labeled Anomaly Ratios. Specifically, as the Labeled Anomaly Ratios decrease, our method’s performance degrades more slowly and maintains relatively high detection performance, while other methods exhibit a significant decline in most datasets. This indicates that our method does not overly rely on the small fraction of labeled anomalies and effectively utilizes the modeling of multiple normal prototypes in a large amount of unlabeled data.
On the other hand, from the perspective of anomaly contamination, as the proportion of labeled anomalies decreases, the proportion of unlabeled anomalies contaminating the unlabeled data increases. Fig. 3 shows that, unlike other methods, our method’s detection performance declines more slowly as the anomaly contamination ratio increases, maintaining a relatively good detection performance. This demonstrates the robustness of our model against anomaly contamination. This robustness is attributed to the model’s design, which estimates the likelihood of each unlabeled sample being normal and uses this estimation as a weight to reduce the impact of contaminating anomalies.
TABLE IV: Ablation study results on 15 datasets. Each column represents the removal of a specific component from the full model. The average AUC-PR and P-values are reported to demonstrate the contribution of each component.
Data Full Model w/o Pretraining w/o $\mathcal{L_{\text{recon}}}$ w/o Decoder w/o $w_{i}$ w/o MNP
DoS 0.946±0.004 0.938±0.014 0.943±0.005 0.942±0.004 0.945±0.003 0.918±0.019
Rec 0.851±0.008 0.840±0.015 0.852±0.012 0.855±0.014 0.836±0.022 0.816±0.014
Bac 0.870±0.005 0.842±0.045 0.847±0.009 0.833±0.039 0.865±0.013 0.802±0.066
Ana 0.910±0.008 0.885±0.048 0.864±0.065 0.855±0.074 0.830±0.064 0.805±0.015
C&C 0.519±0.011 0.488±0.044 0.465±0.061 0.506±0.019 0.498±0.046 0.476±0.036
Attack 0.964±0.012 0.894±0.089 0.952±0.004 0.959±0.015 0.963±0.011 0.875±0.030
DDoS 0.252±0.202 0.161±0.054 0.206±0.170 0.243±0.291 0.242±0.193 0.128±0.023
Okiru 0.518±0.003 0.425±0.190 0.457±0.122 0.520±0.000 0.314±0.220 0.520±0.000
Fraud 0.679±0.021 0.663±0.063 0.649±0.027 0.611±0.082 0.665±0.017 0.601±0.012
VC 0.697±0.005 0.680±0.006 0.689±0.009 0.697±0.008 0.699±0.003 0.633±0.007
VI 0.106±0.012 0.087±0.007 0.099±0.004 0.098±0.010 0.104±0.010 0.088±0.008
Gamma 0.780±0.014 0.770±0.007 0.767±0.016 0.761±0.029 0.768±0.013 0.680±0.016
Census 0.392±0.023 0.380±0.012 0.389±0.016 0.371±0.009 0.374±0.016 0.341±0.017
Pendigits 0.856±0.042 0.844±0.055 0.829±0.070 0.845±0.061 0.821±0.026 0.803±0.067
Bank 0.286±0.010 0.252±0.038 0.227±0.020 0.232±0.014 0.277±0.014 0.231±0.020
Average 0.642 ± 0.025 0.610 ± 0.046 0.616 ± 0.041 0.622 ± 0.045 0.614 ± 0.045 0.581 ± 0.023
P-value - $<$0.0001 0.0001 0.0017 0.0034 $<$0.0001
IV-D Detection Performance on Unseen Anomalies
@@
A crucial aspect of an effective anomaly detection system is its ability to generalize from known anomalies to detect previously unseen ones. This capability is essential for maintaining robustness in dynamic environments where new types of anomalies may emerge. To evaluate our method’s performance on unseen anomalies, we constructed several experimental datasets using the UNSW-NB15 dataset, which contains four types of anomalies: DoS, Rec, Ana, and Bac. In each experiment, two types of anomalies were selected: one used exclusively in the test set (unseen) and the other included in the training and validation sets (seen). Consistent with our weakly supervised anomaly detection settings, the labeled anomaly ratio in the training set was maintained at 1.0%. This setup resulted in 12 different experimental datasets.
@@
The results are presented in Table III, showing the AUC-PR scores for different methods across various combinations of seen and unseen anomaly types. Our method achieved the highest average AUC-PR of 0.765, significantly outperforming other methods by more than 20%. The statistical confidence level for this improvement exceeds 99.0%. This experiment demonstrates that our method can effectively generalize from known anomalies to detect previously unseen ones, maintaining strong detection performance and robustness in dynamic environments. The success of our approach may be attributed to its ability to maintain multiple normal prototypes, which better represent the diverse nature of normal behavior. This focus on distinguishing data points based on their contrast with the prototypes, rather than relying exclusively on previously limited labeled anomalies, likely contributes to the observed improvements.
@@
IV-E Ablation Study
We conducted an ablation study to evaluate the contribution of each component in our model. Table IV summarizes the results across 15 datasets. The full model achieves the highest average AUC-PR of $0.642$, demonstrating the effectiveness of the complete framework. Below, we analyze the impact of removing each component in order.
w/o Pretraining: Removing the pretraining step generally decreases performance, as it negatively impacts the model initialization. Overall, the effect of pretraining is more pronounced in smaller datasets, while its impact diminishes as dataset size increases. For instance, in the relatively small Aposemat IoT-23 datasets such as DDoS ($0.252\pm 0.202\to 0.161\pm 0.054$) and Okiru ($0.518\pm 0.003\to 0.425\pm 0.190$), pretraining significantly boosts performance. In contrast, larger datasets such as UNSW-NB15 DoS ($0.946\pm 0.004\to 0.938\pm 0.014$) show a relatively smaller degradation when pretraining is removed. These results indicate that pretraining provides essential initialization benefits, particularly for datasets with limited size. Although the degradation is moderate, pretraining proves beneficial across diverse scenarios, enhancing the robustness of the model.
w/o $\mathcal{L_{\text{recon}}}$: The reconstruction loss $\mathcal{L_{\text{recon}}}$ is crucial for learning representations of normal data. Removing this component results in significant performance degradation in datasets like Ana ($0.910\pm 0.008\to 0.864\pm 0.065$) and C&C ($0.519\pm 0.011\to 0.465\pm 0.061$). These results demonstrate that reconstruction loss helps model normal patterns and separate them from anomalies effectively. However, its absence has a smaller effect on datasets like Rec, where the AUC-PR remains stable ($0.851\pm 0.008\to 0.852\pm 0.012$).
(a) Synthetic Multi-Modal Data.
(b) Multi-normal Prototypes Anomaly Detection.
(c) Single-normal Prototypes Anomaly Detection.
Figure 5: Anomaly detection results on synthetic data. (a) The synthetic dataset comprises three Gaussian clusters representing normal samples and scattered anomalies. (b) Anomaly detection using multi-normal prototypes successfully captures the three distinct normal modes, assigning low anomaly scores to normal samples while effectively identifying anomalies. (c) Anomaly detection using a single normal prototype fails to adapt to the multi-modal nature of the normal data, resulting in misclassified normal samples and missed anomalies.
w/o Decoder: Removing the decoder (and thus the reconstruction error) shows varied effects. In the Okiru dataset, the AUC-PR remains unchanged ($0.518\pm 0.003$), suggesting that the decoder plays a less significant role for simpler data structures. However, on the Fraud dataset, the AUC-PR drops substantially ($0.679\pm 0.021\to 0.611\pm 0.082$), highlighting its importance for datasets with highly imbalanced distributions.
w/o $w_{i}$: The sample weighting mechanism $w_{i}$ is designed to dynamically adjust the contributions of samples, particularly to address noise, imbalance, and ambiguity in the training process. Its removal results in varying levels of degradation depending on the dataset characteristics. For example, in the Okiru dataset ($0.518\pm 0.003\to 0.314\pm 0.220$), the absence of $w_{i}$ leads to a substantial decline, highlighting its role in filtering noisy or ambiguous samples. Even in datasets with moderate imbalance, such as Rec ($0.851\pm 0.008\to 0.836\pm 0.022$), removing $w_{i}$ reduces performance, demonstrating its utility in enhancing the model’s robustness.
w/o MNP: The exclusion of multi-normal prototypes (MNP) leads to the most significant performance degradation across almost all datasets, confirming its critical role in the framework. Notably, the Attack ($0.964\pm 0.012\to 0.875\pm 0.030$) and DDoS ($0.252\pm 0.202\to 0.128\pm 0.023$) datasets exhibit the largest drops, further demonstrating that single-prototype representations fail to capture the multi-modal nature of normal data.
IV-F Impact of Normal Prototypes: Empirical and Synthetic Analysis
In this section, we examine the effect of varying the number of normal prototypes through both empirical studies on real-world datasets and controlled experiments on synthetic data.
@@ IV-F Impact of Normal Prototypes: Empirical and Synthetic Analysis
As illustrated in Fig. 4, which presents results on four representative real-world datasets, the number of prototypes significantly influences the anomaly detection performance. For most datasets, using multiple prototypes yields better results, supporting our hypothesis that normal data often exhibits multi-modal characteristics. In the Rec dataset, the prototype count has minimal effect, suggesting lower sensitivity to this parameter. However, in the C&C and Census datasets, the number of prototypes plays a critical role—both too many and too few prototypes lead to substantial performance degradation. In the Bank dataset, despite fluctuations in performance with varying prototypes, the results consistently surpass those achieved with a single prototype.
To further validate the multi-modal nature of normal data and the effectiveness of multiple prototypes, we conducted experiments on a synthetic dataset designed with explicit multi-modal characteristics. The synthetic dataset consists of three Gaussian clusters representing normal samples and a small number of uniformly distributed anomalies, as shown in Fig. 5 (a). We compare the anomaly detection results using multi-normal prototypes and a single normal prototype. Fig. 5 (b) demonstrates the effectiveness of multi-normal prototypes in accurately capturing the three distinct modes of the normal data distribution. The model assigns low anomaly scores to normal samples within their respective clusters while effectively identifying anomalies scattered across the feature space. In contrast, Fig. 5 (c) shows the results with a single normal prototype. Here, the model fails to adapt to the multi-modal distribution of the normal data, resulting in higher anomaly scores even for normal samples located farther from the single prototype. Additionally, anomalies positioned in the central region, equidistant from the three normal modes, are often missed due to their proximity to the single prototype. This failure illustrates that a single prototype cannot adequately represent diverse normal patterns, leading to poorer anomaly detection performance.
Figure 6: AUC-PR of our method w.r.t. Three key hyperparameters
IV-G Sensitivity Analysis
We assess the model’s sensitivity to three key hyperparameters: the latent representation dimension $H$, the margin $m_{1}$, and the scaling parameter $\beta$. As shown in Fig. 6, we analyze the trends by selecting four representative datasets (Rec, C&C, Census, and Bank).
For $H$, the model shows relative stability across different values, with $H=8$ providing a good trade-off between performance and computational complexity. Increasing $H$ beyond 8 does not result in significant performance gains and may introduce unnecessary computational overhead. The consistent performance across $H$ values highlights the model’s ability to effectively capture the underlying data structure even with relatively small latent dimensions.
For $m_{1}$, a margin value of approximately $0.02$ achieves the best performance across all datasets. Smaller margins, such as $m_{1}=0.01$, fail to adequately separate the reconstruction losses of normal and anomalous samples, leading to reduced detection accuracy. On the other hand, larger margins, particularly $m_{1}=1.0$ and $m_{1}=10.0$, hinder the model’s capacity to learn anomaly-specific latent representations, resulting in substantial performance degradation.
For $\beta$, which controls the scaling of dynamic weights, the optimal value is around $\beta=1$. Smaller values (e.g., $\beta=0.1$) reduce the effect of dynamic weighting, limiting the model’s ability to suppress noisy samples, while larger values (e.g., $\beta=10$) overemphasize specific samples, potentially causing overfitting or instability. The improvements observed around $\beta=1$ demonstrate its critical role in balancing sample contributions effectively during training.
V Conclusions